Tyrion:

Solve the equation the square root of the quantity x plus 4 minus 3 equals 1 for the variable. Show each step of your solution process.

NoProbBob:

ok so √X+4-3=1

NoProbBob:

Can You solve from there?

Tyrion:

nope

NoProbBob:

Ight lemme break it down foe you

NoProbBob:

ok u got a calculator?

snowflake0531:

\[\sqrt(x+4)-3 = 1\] Can you first add 3 to each side

NoProbBob:

broo i was helpin first.

snowflake0531:

What do you need a calcalator for

NoProbBob:

da square root

snowflake0531:

So you're wrong e.e

snowflake0531:

You have to square both sides you don't need a square root calculator

darkknight:

snowflake0531 wrote:
\[\sqrt(x+4)-3 = 1\] Can you first add 3 to each side
\[\sqrt{x+4}-3=1\]

Tyrion:

ok ill do it myself,,less annoying

NoProbBob:

not a square root calualor, jus a reg calc homie

snowflake0531:

@noprobbob wrote:
not a square root calualor, jus a reg calc homie
My meaning was.. a calculator

NoProbBob:

U got dis bro. All u gotta do is add em up

darkknight:

darkknight wrote:
snowflake0531 wrote:
\[\sqrt(x+4)-3 = 1\] Can you first add 3 to each side
\[\sqrt{x+4}-3=1\]
add 3 to both sides, square both sides then subtract both sides by 4. Boom u got x

Tyrion:

yall kan stop responding now thank you :)